OpenStudy (anonymous):

tan theta=12/5, sin theta <0 how would i find the rest of the trig values?

Parth (parthkohli):

Are you aware that tan(theta) = sin(theta)/cos(theta)?

OpenStudy (anonymous):

yes i am

OpenStudy (anonymous):

i understand that the top part of sin is 12, but when i checked in the back up the book the answer was sin= -12/13 and i don't understand how the -13 came into play. I know it has to do with a^2+b^2=c^2 with 5^2+12^2=13^2 but i don't fully understand it.

Parth (parthkohli):

Hey, I am back

Parth (parthkohli):

It is given that sin(theta) is negative and tan(theta) is positive. Because tan = sin/cos, it is clear that cos(theta) is also negative.

Parth (parthkohli):

The sides of a triangle can be negative as well. The concept me may be confusing, but yeah. Have you heard of the unit circle?